OpenStudy (anonymous):

A cylindrical water tank with 14m Diameter and 13m height. What is the capacity of tank in terms of Litters ??

OpenStudy (zzr0ck3r):

\(V=\pi r^2h\)

OpenStudy (anonymous):

What is the capacity of tank in terms of Litters ?

OpenStudy (zzr0ck3r):

Turn the meters to cm

OpenStudy (anonymous):

convert meter to cm @sampathk

OpenStudy (zzr0ck3r):

so \(V=\pi*650^2*700 \ cm^3\)

OpenStudy (anonymous):

no @Jhannybean

OpenStudy (anonymous):

i thought height is 13m @zzr0ck3r

OpenStudy (zzr0ck3r):

oh im out of it....

OpenStudy (zzr0ck3r):

square it

OpenStudy (zzr0ck3r):

\(V=\pi*700^2*1300 \ cm^3 \)

OpenStudy (zzr0ck3r):

Together we can get this:P

OpenStudy (anonymous):

solve it @sampathk

OpenStudy (anonymous):

It Is a volume =2001194520 cm^3 .. How will you express it in-terms of Litters ??

OpenStudy (anonymous):